OpenStudy (anonymous):

What is the solution of 5d+6-3d=12

OpenStudy (ja1):

Can you combine the like terms?

OpenStudy (anonymous):

yes

OpenStudy (ja1):

Ok so what do you get after you combine them?

OpenStudy (anonymous):

8d?

OpenStudy (ja1):

No, see you always need to remember to carry the sign, so you would do 5d - 3d

OpenStudy (anonymous):

so its 2d?

OpenStudy (anonymous):

will the answer be 3?

OpenStudy (texaschic101):

5d + 6 - 3d = 12 2d + 6 = 12 --- subtract 6 from both sides 2d = 12 - 6 2d = 6 -- divide by 2 d = 6/2 d = 3 yes...the answer is 3

OpenStudy (ja1):

Yes, sorry I was taking care of something but I see texas has it under control :)